OpenStudy (anonymous):

Solve the formula for h. What is the height of a cylinder when the surface area is 75.36 square inches and the radius is 2 inches? Use 3.14 for pi. (Formula below)

OpenStudy (anonymous):

\[S = 2 \pi rh + 2 \pi r ^{2}\]

OpenStudy (anonymous):

@UnkleRhaukus can you help me?

OpenStudy (anonymous):

@snark can you help me?

OpenStudy (anonymous):

So you just substitute in all the numbers!

OpenStudy (anonymous):

that's the problem. the question didn't give me any numbers.

OpenStudy (anonymous):

75.36 = 2 * 3.14 * 2 * h + 2 * 3.14 * r^2

OpenStudy (anonymous):

whoops, 75.36 = 2 * 3.14 * 2 * h + 2 * 3.14 * 2^2

OpenStudy (anonymous):

75.36 = 12.56h + 25.12 then do some algebra 50.24 = 12.56h

OpenStudy (anonymous):

and divide both sides by 12.56 h = 4

OpenStudy (anonymous):

did i every tell you cholo71796, that you are a GENIUS!!

OpenStudy (anonymous):

Remember that you should be able to solve for an answer if you know all the variables but one and you know how all of them relate to each other. In this equation, the variable are the surface area, radius and the height. Since they give us the surface area and radius, we can solve for height!

OpenStudy (anonymous):

Thanks!